Prove ∫∞0 𝑒 βˆ’π‘₯ 2 π‘₯ 2 𝑑π‘₯ = βˆšπœ‹/ 4


1
Expert's answer
2021-11-10T17:11:40-0500

Β 

"\int \mathrm{e}^{-x^{2}}x^{2} \mathrm{~d} x=" "\int x^{2} \mathrm{e}^{-x^{2}} \mathrm{~d} x"

Using Integration by parts

Β "=-\frac{x \mathrm{e}^{-x^{2}}}{2}-\int-\frac{\mathrm{e}^{-x^{2}}}{2} \mathrm{~d} x" ...(1)

Now solving:

"\int-\frac{\mathrm{e}^{-x^{2}}}{2} \mathrm{~d} x"

Apply linearity:

Β "\begin{aligned}

&=-\frac{\sqrt{\pi}}{4} \int \frac{2 \mathrm{e}^{-x^{2}}}{\sqrt{\pi}} \mathrm{d} x \\

&=-\frac{\sqrt{\pi} \operatorname{erf}(x)}{4}

\end{aligned}"

Put the above value in eqn .(1)

Β "\begin{aligned}

&-\frac{x \mathrm{e}^{-x^{2}}}{2}-\int-\frac{\mathrm{e}^{-x^{2}}}{2} \mathrm{~d} x \\

&=\frac{\sqrt{\pi} \operatorname{erf}(x)}{4}-\frac{x \mathrm{e}^{-x^{2}}}{2}+C

\end{aligned}"

So, "\int x^{2} \mathrm{e}^{-x^{2}} \mathrm{~d} x=\frac{\sqrt{\pi} \operatorname{erf}(x)}{4}-\frac{x \mathrm{e}^{-x^{2}}}{2}+C"

And "\int_{0}^{\infty} x^{2} \mathrm{e}^{-x^{2}} \mathrm{~d} x=[\frac{\sqrt{\pi} \operatorname{erf}(x)}{4}-\frac{x \mathrm{e}^{-x^{2}}}{2}+C]_{0}^{\infty}=\frac{\sqrt{\pi}}{4}"

Β